OpenStudy (anonymous):

Estimate the answer by rounding. 932 - 676

OpenStudy (anonymous):

So: 932 - 676 = 256 <----- this is the exact answer ------------------------------------------------------------ so then: (if rounding to the nearest hundred) 932 rounds to approximately to 900 676 rounds to approximately to 700 ========================================== so then: 900 - 700 = 200 ++++++++++++++++++++++++++++++++++++++++++++ Or: (rounding to the nearest tens) 932 rounds to 930 676 rounds to 680 so then: 930 - 680 = 250